OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form.

OpenStudy (anonymous):

The zeros are: 3, -13, and 5 + 4i The choices were: A) f(x) = x^4 - 8x^3 - 12x^2 + 400x - 1599 B) f(x) = x^4 - 200x^2 + 800x - 1599 C) f(x) = x^4 - 98x^2 + 800x - 1599 D) f(x) = x^4 - 8x^3 + 12x^2 - 400x + 1599

OpenStudy (anonymous):

I've asked this one before, but I still don't understand it.

OpenStudy (mathstudent55):

Each real zero "a", will give you a factor of (x - a). Each complex zero a + bi will also have to have its complex conjugate as a zero, adding two more factors: (x + (a + bi)) and (x - (a - bi)) Then multiply all the factors together to find the polynomial.

OpenStudy (mathstudent55):

In your case, you have two real zeros: 3 and -13. Based on what I wrote above, which factors do you get from those two zeros?

OpenStudy (anonymous):

So I'll have (x-3), (x+13), (x+(5+4i)), and (x-(5+4i)). Right?

OpenStudy (mathstudent55):

Almost. (x - 3) and (x + 13) are correct. You have to be careful with the complex zeros. Since you have complex zero 5 + 4i given, its complex conjugate 5 - 4i is also a zero. Now write the factors for the complex conjugate zeros.

OpenStudy (anonymous):

Oh, so the second one would be (x-(5-4i)) instead?

OpenStudy (anonymous):

I'm confused on the next step. How do I multiply all three of them?

OpenStudy (mathstudent55):

Not quite. The 2 complex roots are: \(x - (5 + 4i)) and (5 - (5 - 4i))

OpenStudy (anonymous):

How did you get the (5-(5-4i))?

OpenStudy (mathstudent55):

The given complex root is 5 + 4i, right?

OpenStudy (loser66):

@mathstudent55 please solve it, I don't get the answer,

OpenStudy (anonymous):

Yesh.

OpenStudy (mathstudent55):

I'm getting there. One step at a time.

OpenStudy (mathstudent55):

That automatically gives us a second complex root of 5 - 4i. The reason for this is that we are told the coefficients of the polynomial are real. A polynomial with real coefficients can have real roots and/or complex roots, but the complex roots must come in complex conjugate pairs.

OpenStudy (anonymous):

Okay. I understand the 5+4i and 5-4i, but not the other half.

OpenStudy (mathstudent55):

Now we know we have complex roots of 5 + 4i and 5 - 4i. All our roots are: 3, -13, 5 + 4i, 5 - 4i. Ok?

OpenStudy (anonymous):

Got it.

OpenStudy (mathstudent55):

Every root "a" generates a factor of the form (x - a). We have 4 factors that need to be multiplied together: \( (x - 3)(x + 13)(x - (5 + 4i)((x - (5 - 4i)) \)

OpenStudy (anonymous):

Those are the ones I said the second time. :p -_-

OpenStudy (mathstudent55):

Now we have to multiply together all those factors. We can start by multiplying the first two, but a shortcut is to make sure you multiply the last two together.

OpenStudy (mathstudent55):

I see it now. I wrote 5 instead of x above. Sorry. It was a typo.

OpenStudy (anonymous):

I thought so. You're good. I understand the roots now. How do I multiply them though?

OpenStudy (mathstudent55):

The last two factors that have the complex roots in them are of the form of a product of a sum and a difference. They follow this form: \( (a + b)(a - b) = a^2 - b^2\)

OpenStudy (mathstudent55):

We can take advantage of this as a shortcut.

OpenStudy (anonymous):

You're talking about the 5+4i and 5-4i, right?

OpenStudy (mathstudent55):

Yes.

OpenStudy (mathstudent55):

We manipulate the factors with the complex roots into a product of a sum and a difference. \( (x - 3)(x + 13)\color{red}{(x - (5 + 4i)((x - (5 - 4i))} \) \(= (x - 3)(x + 13) \color{red}{(x - 5 - 4i)(x - 5 + 4i)} \) \(= (x - 3)(x + 13) \color{red}{((x - 5) - (4i))((x - 5) + (4i))} \)

OpenStudy (anonymous):

Okay... I think I got it so far..

OpenStudy (mathstudent55):

Now we multiply the last two factors together using the product of a sum and a difference.

OpenStudy (mathstudent55):

\(= (x - 3)(x + 13) \color{red}{((x - 5)^2 - (4i)^2)} \) \(= (x - 3)(x + 13) \color{red}{(x^2 - 10x + 25 - (-16) )} \)

OpenStudy (anonymous):

Which parts do we plug into the a^2-b^2? Oh. Nevermind.

OpenStudy (mathstudent55):

\(= (x - 3)(x + 13) (x^2 - 10x + 25 + 16)\) \(= (x - 3)(x + 13) (x^2 - 10x + 41)\)

OpenStudy (anonymous):

So then we need to multiply x^2+10x-39 and x^2-10+41?

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

How do you multiply those.. Is there a thing like FOIL but for three's?

OpenStudy (mathstudent55):

\(= (x^2 +10x - 39) (x^2 - 10x + 41)\) Multiply every term of the first polynomial by every term of the second polynomial. Start by taking the first term of the first polynomial and multiplying by all terms of the second one. Then take the second term of the first polynomial and multiply by every term of the second one. Then do the same with the third term. Finally, combine like terms.

OpenStudy (mathstudent55):

\(= (x^2 +10x - 39) (x^2 - 10x + 41)\) \(= x^4 - 10x^3 + 41x^2 + 10x^3 - 100x^2 + 410x - 39x^2 + 390 x - 1599\)

OpenStudy (mathstudent55):

\(= x^4 - 98x^2 + 800x - 1599\)

OpenStudy (anonymous):

That's what I got, too! :D So it's c. Yay!

OpenStudy (anonymous):

I have another one. You up for it? :p

OpenStudy (anonymous):

the hardest part is finding a quadratic with zero of \( 5 + 4i \) but it is easy if you work backwards \[x=5+4i\\ x-5=4i\\ (x-5)^2=(4i)^2=16\\ x^2-10x+25=-16\\ x^2-10x+41=0\]

OpenStudy (anonymous):

what is even easier is to memorize that if \(a+bi\) is a the root of a quadratic, then it is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (mathstudent55):

@satellite73 Thanks for the insight.

OpenStudy (anonymous):

Okay. :) The other question I have isn't like this one.

OpenStudy (anonymous):

yw

OpenStudy (mathstudent55):

Pls start a new post for the new problem.

OpenStudy (anonymous):

Fine..